An object with a mass of 50 kg is being dragged along a horizontal floor by a horizontal force of 30 N. If the friction force opposing the motion is a steady 20 N, what is the acceleration of the body
1
Expert's answer
2012-11-13T11:48:42-0500
According to the second Newton's law the acceleration of an object is

a = F/m = (30[N]-20[N])/50[kg] = 0.2[m/s²].